### v3

# Neural Diffusion Processes for Physically Interpretable Survival Prediction

Abstract
We introduce DeepFHT, a survival-analysis framework that couples deep neural networks with first hitting time (FHT) distributions from stochastic process theory. Time to event is represented as the first passage of a latent diffusion process to an absorbing boundary. A neural network maps input variables to physically meaningful parameters including initial condition, drift, and diffusion, within a chosen FHT process such as Brownian motion, both with drift and driftless. This yields closed-form survival and hazard functions and captures time-varying risk without assuming proportional-hazards.
We compare DeepFHT with Cox regression using synthetic and real-world datasets. The method achieves predictive accuracy on par with the state-of-the-art approach, while maintaining a physics-based interpretable parameterization that elucidates the relation between input features and risk. This combination of stochastic process theory and deep learning provides a principled avenue for modeling survival phenomena in complex systems.
Suggested keywords
preprint: APS/123-QED
I Introduction
Survival analysis is central in many applications across medicine, engineering, economics and finance. It concerns time-to-event modeling: given a process that can generate an event of interest (e.g., death from disease, failure due to wear), the goal is to estimate the probability that an event occurs at any time $t>0$ for an individual described by some input variables (or features, or covariates). Unlike standard regression settings, survival data are characterized by censoring, which means that for some instances, the exact event time is not observed (for example, when individuals remain event-free at the end of the study), and only the last recorded follow-up time is available.
Traditional approaches to survival modeling rely on strong statistical assumptions linking input variables and risk. The Cox proportional hazards (CoxPH) model [1] remains the most widely used and best established method. The proportional hazards assumption implies that the instantaneous risk of event for two individuals differs by a constant factor over time. The CoxPH model is also linear, making it clear how each single input variable affects the outcome, but at the expense of missing interactions between features. In its original form, this relation is modeled through a linear regression on the features, though many extensions have been developed to relax linearity and improve performance in high-dimensional settings [2, 3, 4].
Despite its success, Cox regression is limited by the proportional hazards (PH) assumption, which is often unrealistic. Consequently, many alternatives have been proposed that dispense with the PH constraint, ranging from classical statistical formulations [5, 6] to modern methods that incorporate machine learning and deep learning techniques [7, 8, 9].
Among these, first hitting time (FHT) models are of particular interest for this work. FHT models assume that event times are distributed as the times of first passage of a latent stochastic process, underlying the event-generating process observed, through an absorbing barrier. This approach overcomes the PH assumption, since FHT distributions for many stochastic processes allow for time-varying hazard ratios between different individuals (e.g., Lévy distribution, Inverse Gaussian distribution).
A branch of survival methods seeks to model the distribution of event times directly, without imposing a parametric process model. Examples include HitBoost [10] and DeepHit [11], which use machine learning and deep neural networks to approximate the event-time density $f(t)$ from data. While these approaches often achieve strong predictive performance, they do so in a purely data-driven manner: the estimated distribution is not tied to parameters of an underlying stochastic process. As a result, the learned models provide limited physical or mechanistic interpretability, and it remains difficult to understand how features influence the latent dynamics that give rise to observed survival outcomes. Conversely, parametric FHT alternatives [12, 13, 14] currently suffer from the opposite drawbacks. These models learn the process parameters on which the distribution of first hitting times depends from data, using simpler regression and machine learning techniques, yielding more physically interpretable results at the expense of model expressiveness.
The present work addresses these complementary limitations by introducing DeepFHT, a framework that unites parametric FHT modeling with deep neural networks. In DeepFHT, a neural network maps input features to the parameters of an underlying stochastic process—such as initial condition, drift, and diffusion—that governs the distribution of first hitting times. To our knowledge, this is the first application of modern deep learning to parametric FHT survival models. The hybrid design combines the representational power of neural networks with the mechanistic grounding of FHT processes, yielding closed-form survival and hazard functions while maintaining a direct physical interpretation of parameters. In this way, DeepFHT provides a principled alternative to black-box survival models, providing predictions that are not only accurate but also anchored in a dynamical description that links covariates to the latent process generating events.
We demonstrate the capability of this framework on synthetic and publicly available clinical survival datasets, comparing predictive power against the CoxPH model. DeepFHT achieves competitive performance across clinical benchmarks and clearly outperforms Cox regression on synthetic data designed to probe the model’s capabilities in scenarios with nonlinearities and non-proportional hazards.
Finally, we assess interpretability by examining the correlations between predicted process parameters, empirical risk, and features. The results show that the similarity between patients outcomes is intuitively encoded as distance in the space of parameters, and that DeepFHT can recover clinically meaningful relationships and has the potential to highlight new correlations in complex, high-dimensional data.
The remainder of the paper is organized as follows. Sections II and III provide theoretical background, reviewing the fundamentals of survival analysis and introducing the relevant aspects of stochastic process theory. Section IV presents the proposed models and the experimental setup, while the results and the physics-based interpretability of the models are discussed in Section V.
II Survival analysis framework
Survival datasets are typically represented as triplets $(\mathbf{x}_{i},\delta_{i},T_{i})$ for $i=1,\dots,n$, where:
  • $n\in\mathbb{N}$ is the dataset size;
  • $\mathbf{x}_{i}\in\mathbb{R}^{m}$ is a vector of $m$ features;
  • $\delta_{i}\in\{0,1\}$ is an indicator that denotes whether the event was observed ($\delta_{i}=1$) or the instance was right censored, i.e., end of observation without occurrence of the event ($\delta_{i}=0$);
  • $T_{i}\in\mathbb{R}_{>0}$ is the time of event occurrence or censoring.
The goal of survival analysis is to predict the probability distribution of event times $f(t)$ and some relevant, related functions:
  • survival function $S(t)=\Pr(T>t)$
  • cumulative distribution function F(t)=1S(t)F(t)=1-S(t)
  • hazard function
    $h(t)=\lim_{\Delta t\to 0^{+}}\frac{\Pr(t\leq T<t+\Delta t\mid T\geq t)}{\Delta t}=\frac{f(t)}{S(t)},$
The survival function describes the probability of surviving beyond time $t$, while the hazard captures the instantaneous risk of event occurrence, given survival up to $t$. Estimation typically focuses on conditional forms $S_{\bm{\theta}}(t|\mathbf{x})$ and $h_{\bm{\theta}}(t|\mathbf{x})$, which quantify the dependence of survival dynamics on the features, through a set of parameters $\bm{\theta}$. If one assumes independent samples and non-informative censoring, the parameters can be determined with a maximum likelihood estimation (MLE) from the data [15].
A widely adopted semi-parametric approach for modeling how covariates affect the time-to-event distribution is the Cox Proportional Hazards model [1]. It assumes that the hazard rate for an individual with features vector $\mathbf{x}$ is proportional to a baseline hazard function $h_{0}(t)$ and can be factorized as:
$h(t\mid\mathbf{x})=h_{0}(t)\exp\left\{\bm{\beta}^{\top}\mathbf{x}\right\},\quad S(t\mid\mathbf{x})=\left[S_{0}(t)\right]^{\exp\left\{\bm{\beta}^{\top}\mathbf{x}\right\}}.$ (1)
This assumption implies that the ratio of hazards for any two individuals is constant over time, depending only on their covariates and not on $t$ itself. Unlike fully parametric methods, CoxPH does not require specifying a functional form for the baseline hazard during the MLE, since the time-dependent components cancel out in the likelihood.
III First hitting time models for survival analysis
A natural probabilistic representation of survival processes is provided by the theory of first hitting times. Consider a one-dimensional continuous-time stochastic process $\{X(t)\}_{t\geq 0}$, with $X(t)\in\mathbb{R}$, described by a probability distribution $p(x,t)$. Let $\Gamma=\{x\leq x_{b}\}\subset\mathbb{R}$ be a target set and $X(t=0)=x_{0}>x_{b}$ the initial condition. In this setup, $x_{b}$ can be considered as a barrier between the target set and the initial condition and this is how it will be referred to in the rest of the paper. The first hitting time is then
$T=\inf\{t>0:X(t)\in\Gamma\},$ (2)
that is, the earliest time at which the process crosses the barrier in $x_{b}$. When the barrier is absorbing (i.e., killing the process upon transition), the distribution $f(T)$ of first hitting times coincides with the event time distribution in survival analysis. This condition is easily imposed by the Dirichlet condition $p(x_{b},t\mid x_{0})=0$. Without losing generalizability, we will always consider positive initial conditions $x_{0}>0$ and an absorbing barrier in $x_{b}=0$. This amounts to a generic choice for a process in $X(t)\in\mathbb{R}$.
Since all relevant survival functions can be derived from the distribution of first hitting times $f(T)$, it is crucial to have analytical solutions for this quantity. In this work, we consider two stochastic processes whose theory is well developed and for which an analytical solution exists: the Brownian motion and the arithmetic Brownian motion (Brownian motion with drift). For a more complete and detailed discussion of FHT theory and models, including but not limited to their application to survival analysis, we refer to [16, 12, 17].
III.1 First hitting time and survival distributions for Brownian motion
Brownian motion is a pure diffusion process described by a probability distribution $p(x,t\mid x_{0})$ that evolves according to the Fokker–Planck equation:
$\frac{\partial}{\partial t}p(x,t\mid x_{0})=D\frac{\partial^{2}}{\partial x^{2}}p(x,t\mid x_{0}),$ (3)
with constant diffusion coefficient $D>0$ and initial condition $p(x,0\mid x_{0})=\delta(x-x_{0})$. Imposing an absorbing barrier at $x_{b}=0$ corresponds to the Dirichlet boundary condition
$p(0,t\mid x_{0})=0,\qquad x_{0}>0.$
The method of images yields the transition probability density (exact solution):
$\small p(x,t\mid x_{0})=\frac{1}{\sqrt{4\pi Dt}}\left[\exp\!\left(-\frac{(x-x_{0})^{2}}{4Dt}\right)-\exp\!\left(-\frac{(x+x_{0})^{2}}{4Dt}\right)\right]$ (4)
This solution describes a Brownian motion starting at $x_{0}>0$ with an absorbing barrier at the origin.
From the probability distribution (4), all relevant survival quantities can be derived. We focus specifically on:
  • Survival function
    $S(t)=\int_{0}^{\infty}p(x,t\mid x_{0})\,dx=\operatorname{erf}\!\left(\frac{x_{0}}{\sqrt{4Dt}}\right),$ (5)
  • Failure density
    $f(t)=-\frac{d}{dt}S(t)=\frac{x_{0}}{2\sqrt{\pi}}\,\frac{\exp\!\big(-x_{0}^{2}/(4Dt)\big)}{(Dt)^{3/2}},$ (6)
    the distribution of first hitting times (i.e., event times).
The Brownian FHT model is characterized by a failure density that has the form of a Lévy distribution, depending only on process parameters $x_{0},D$.
III.2 First hitting time and survival distributions for arithmetic Brownian motion
Adding a constant drift term $\mu$ to Brownian motion yields the arithmetic Brownian motion. For this process, the Fokker-Planck equation becomes:
$\frac{\partial}{\partial t}p(x,t\mid x_{0})=-\,\mu\,\frac{\partial}{\partial x}p(x,t\mid x_{0})+D\frac{\partial^{2}}{\partial x^{2}}p(x,t\mid x_{0}),$ (7)
with a constant diffusion coefficient $D>0$. For an initial condition $p(x,0\mid x_{0})=\delta(x-x_{0})$, the free space solution is Gaussian with the mean shifted linearly in time:
$p(x,t\mid x_{0})=\frac{1}{\sqrt{4\pi Dt}}\exp\!\left(-\frac{(x-x_{0}-\mu t)^{2}}{4Dt}\right).$ (8)
When introducing an absorbing barrier at $x_{b}=0$, it becomes, via the method of images:
$p(x,t\mid x_{0})$ $=\frac{1}{\sqrt{4\pi Dt}}\Bigg[\exp\!\left(-\frac{(x-x_{0}-\mu t)^{2}}{4Dt}\right)$ (9)
$\qquad-\exp\!\left(-\frac{\mu x_{0}}{D}\right)\,\exp\!\left(-\frac{(x+x_{0}-\mu t)^{2}}{4Dt}\right)\Bigg],$
valid for $x>0$, $x_{0}>0$.
From (9), the survival analysis quantities follow:
  • Survival function
    $S(t)$ $=\int_{0}^{\infty}p(x,t\mid x_{0})\,dx$ (10)
    $=\Phi\!\left(\frac{x_{0}+\mu t}{\sqrt{2Dt}}\right)-\exp\!\left(-\frac{\mu x_{0}}{D}\right)\,\Phi\!\left(\frac{\mu t-x_{0}}{\sqrt{2Dt}}\right),$
    where $\Phi(\cdot)$ denotes the standard normal cdf;
  • Failure density
    $f(t)=\frac{x_{0}}{\sqrt{4\pi Dt^{3}}}\,\exp\!\left(-\frac{(x_{0}+\mu t)^{2}}{4Dt}\right),$ (11)
    which corresponds to an inverse Gaussian distribution.
In the limit $\mu\to 0$, these expressions reduce to the driftless Brownian case, with the failure density converging to the Lévy distribution. Given sufficient time, failure is certain only for $\mu<0$.
IV Methods
IV.1 DeepFHT
As outlined in Sec. III, survival outcomes can be modeled by first hitting time (FHT) distributions with absorbing barriers. The corresponding survival functions and event-time distributions are determined by the choice of underlying process, depending explicitly on its parameters.
Figure 1: Example output of the Lévy FHT model. Individual-specific survival functions are computed from the neural network–predicted parameters.
To model these functions and distributions from data, we employ a feedforward neural network $f_{\theta}$ that maps features $\mathbf{x}_{i}\in\mathbb{R}^{m}$ to a vector of distributional parameters $\mathbf{p}_{i}=f_{\theta}(\mathbf{x}_{i})\in\mathbb{R}^{d}$, i.e. the parameters of the underlying stochastic process. The parameterizations $\mathbf{p}_{i}$ produced by the network are then used to evaluate survival quantities such as $S(t\mid\mathbf{x}_{i})=S(t;\mathbf{p}_{i}),F(t\mid\mathbf{x}_{i})=F(t;\mathbf{p}_{i})$, where $S$ and $F$ denote the functional forms of the survival and failure distributions of the chosen FHT model.
The dimension $d$ of the output vector $\mathbf{p}$ depends on the parametrization choice. As anticipated in Section III, in this paper we will focus on Brownian and arithmetic Brownian motion, identified by their FHT distributions 6,11, with the following parameterizations:
  • Lévy distribution: $\mathbf{p}=\big(x_{0},D\big)_{\theta}$,
  • Inverse Gaussian distribution: $\mathbf{p}=\big(x_{0},\mu\big)_{\theta}$ with $\mu<0$, $D=1$,
For the inverse–Gaussian model we fix $D=1$ without loss of generality. The FHT law depends on $\mu_{T}=x_{0}/|\mu|$ and $\lambda=x_{0}^{2}/(2D)$, thus $D$ only sets the time scale and can be absorbed by rescaling time (or $\mu$). We therefore parameterize the model by $(x_{0},\mu)$ with $\mu<0$.
Notably, both parameterizations above allow for non-proportional hazards, reflecting the flexibility of FHT-based models to capture complex time dependencies in risk. Figure 1 illustrates a representative model output, where survival functions are computed from the estimated distributional parameters for a set of individuals. It is also worth noting that this model can easily be extended to any other probability distribution of choice. It is sufficient to specify the functional form and its parameterization and to modify the size of the output layer accordingly.
Physics-based interpretability
A distinctive feature of DeepFHT is its physics-based interpretability. Because the model is parametrized by quantities with direct physical meaning (initial condition, drift, and diffusion) each individual is represented as a point in a latent parameter space whose geometry reflects the underlying process dynamics. The influence of every parameter on event-time distribution is analytically known, allowing regions of higher or lower risk to be distinguished and trajectories of individuals to be compared qualitatively in this space. In the specific Lévy (6) and inverse Gaussian (11) cases, the resulting two-dimensional parameterization makes these relationships visually intuitive and enables quantitative notions of similarity (e.g., Euclidean distance) between subjects.
To visualize risk in parameter space, we interpolate event times from the set of uncensored training instances $U=\{\mathbf{p}_{i}\mid\delta_{i}=1\}$ via inverse-distance weighting:
$T(\mathbf{p})=\frac{\sum_{\mathbf{p}_{i}\in U}w_{i}(\mathbf{p})\,T_{i}}{\sum_{\mathbf{p}_{i}\in U}w_{i}(\mathbf{p})},\qquad w_{i}(\mathbf{p})=\frac{1}{\sqrt{d(\mathbf{p},\mathbf{p}_{i})}}$ (12)
Here $d(\cdot,\cdot)$ is a distance in the model parameter space. The resulting $T(\mathbf{p})$ acts as a proxy for risk and supports distance-based similarity analyses.
IV.2 Loss function: Brier loss
Training is performed by minimizing a custom Brier loss, adapted from the Brier score [18, 19] (see Sec. IV.3) commonly used to evaluate probabilistic forecasts. The loss is defined as
$\mathcal{L}=\sum_{t\in U}\sum_{i}\Big[$ $I(T_{i}\leq t\wedge\delta_{i}=1)\,(S(t|\mathbf{x}_{i}))^{2}$ (13)
$+I(T_{i}>t)\,(1-S(t|\mathbf{x}_{i}))^{2}\Big],$
where $U=\{T_{j}:\delta_{j}=1\}$ is the set of unique event times where predictions are evaluated. $S(t|\mathbf{x}_{i})$ denotes the survival probability predicted for subject $i$, and $I(\cdot)$ is the indicator function.
IV.3 Experimental setup
The experimental evaluation of DeepFHT tests model performance when modeling first hitting time distributions for diffusion processes (Lévy, Inverse Gaussian). As an external benchmark, we also include the Cox proportional hazards model, the standard reference in survival analysis.
The procedure consists of two phases. First, for each candidate distribution, we perform 5-fold cross-validation on a training set comprising $80\%$ of the available data, with the remaining $20\%$ held out for testing. Each split preserves the original ratio of observed to censored instances. Cross-validation is used to select the optimal architecture and hyperparameters (see Appendix A), based on average validation performance measured over 100 trials. For hyperparameter search, we use the Optuna framework [20], which implements efficient search strategy for exploring the hyperparameter space.
After hyperparameter optimization, models are retrained on the full training set and evaluated on the held-out test set. For each performance metric, mean and standard deviation are estimated via bootstrap with resampling: 100 datasets of equal size are drawn with replacement from the test set, each preserving the censoring ratio.
Performance metrics
Antolini’s C-index.  We use the time-dependent concordance index (also referred to as Antolini’s C-index) [21] as the performance metric of choice, for both validation and testing. It is a rank statistic that measures agreement between predicted risks and observed survival times as the probability that, among two comparable individuals, the one experiencing the event earlier is assigned a higher predicted risk (or equivalently a lower survival probability). Antolini’s C-index generalizes the original Harrel’s C-index [22] to models where the predicted risk ranking between individuals is not fixed (like in CoxPH) but changes at different times.
$\small c=\frac{\sum_{(i,j)}w(T_{i})I(T_{i}<T_{j}\land\delta_{i}=1\land S(T_{i}\mid\mathbf{X}_{i})<S(T_{i}\mid\mathbf{X}_{j}))}{\sum_{(i,j)}w(T_{i})I(T_{i}<T_{j}\land\delta_{i}=1)}.$ (14)
where $w(T_{i})$ are inverse probability of censoring weights (IPCW) [23], used to correct for bias introduced by censoring. This time-dependent version is a censoring-aware equivalent to the ROC-AUC metrics and it is more appropriate for evaluating models whose predictions vary with time.
Integrated Brier score.  The Brier score [18] measures predictive accuracy as the mean squared difference between predicted survival probabilities and observed outcomes $\delta_{i}$. Its time-dependent extension [19] evaluates survival probabilities $S(t\mid\mathbf{X}_{i})$ at arbitrary times $t$ and uses IPCW. This yields a curve $B(t)$ describing calibration and discrimination across time. To summarize performance over the full observation window, we report the integrated Brier score (IBS),
$B=\frac{1}{t_{\max}-t_{0}}\int_{t_{0}}^{t_{\max}}B(t)\,\mathrm{d}t,$ (15)
where lower values indicate better overall predictive accuracy and a Brier score of 0.25 corresponds to a completely uninformative model. In our experiments, the IBS is used as an additional metric during the testing phase only.
Datasets
We evaluate DeepFHT on publicly available clinical survival datasets, accessed through the SurvSet [24] repository, as well as on a synthetic dataset designed to test non-proportional hazards and nonlinear relationships between features and risk.
All datasets undergo identical preprocessing: categorical features are one-hot encoded; missing values are imputed using the mean or median for numerical variables (depending on skewness) and the mode for categorical variables; finally, all features are scaled.
GBSG2:  Data from the German Breast Cancer Study Group 2 [25], comprising 686 patients with node-positive primary breast cancer. Eight clinical features are available, with $56\%$ censoring.
Framingham:  Derived from the Framingham Heart Study [26], including 4 699 participants and 7 cardiovascular health features.
SUPPORT2:  A multi-center observational study of seriously ill hospitalized adults in the United States [27]. The SurvSet version contains 9 105 records with 35 features; censoring accounts for $32\%$.
NonPH:  A synthetic dataset introduced in [28], specifically designed to violate proportional hazards and enforce nonlinear feature–hazard relationships. It contains 10 000 individuals with 20 normally distributed features. Event times are generated by partitioning the observation period into 16 intervals and assigning individual-specific failure densities:
$p_{i}=\frac{\exp(16x_{i})}{\sum_{j=1}^{16}\exp(16x_{j})},$
where $x_{i}$ are features. This construction yields non-proportional hazards by design (see Appendix B for more details).
V Results
Figure 2: Performance across clinical and synthetic datasets. Scatterplots with error bars for C-index (↑, brown, top row) and IBS (↓, dark blue, bottom row) across datasets (GBSG2, Framingham, SUPPORT2, NonPH). For each dataset, three models are shown: Lévy, inverse Gaussian, and CoxPH. Mean values and standard deviations are computed via 100 bootstrap iterations.
V.1 Model performance
Predictive test results are summarized in Fig. 2. Across datasets, models based on the more physically interpretable Lévy and inverse Gaussian distributions demonstrates competitive performance, consistently matching or surpassing Cox regression in terms of Antolini’s C-index, exhibiting strong robustness and stability at test time.
On the synthetic NonPH dataset, DeepFHT significantly outperforms CoxPH with all tested distributions, underscoring the gains achievable in scenarios with nonlinear effects and violations of the proportional hazards assumption.
Within the clinical datasets, the comparable performance of DeepFHT and CoxPH on GBSG2 and Framingham suggests that in these cases the assumptions of linearity and proportional hazards may not be too restrictive. This points instead to potential limitations in the expressiveness or complexity of the datasets themselves. By contrast, on the higher-dimensional SUPPORT2 dataset, the inverse Gaussian variant achieves a clear improvement over Cox regression, illustrating the advantage of relaxing linear and proportional hazards assumptions in more complex settings.
V.2 Empirical analysis of parameter-space interpretability
Building upon the interpretability introduced in Section IV.1, we now examine how the learned parameters encode survival dynamics in practice. Specifically, we visualize event times and clinical features in the two-dimensional parameter spaces of the Lévy and inverse Gaussian models.
Event times in parameter space.  Using the IDW interpolation defined in Section IV.1, we map each point in the parameter spaces to an estimated event time $T(\mathbf{p})$, computed from uncensored training instances. These maps serve as risk proxies and allow to test whether the model encodes similarity as proximity in parameter space.
Figure 3: Event times in the parameter spaces of Deep FHT models. Left: Framingham dataset in the space of Lévy Deep FHT ($\{x_{0},D\}$). Right: Framingham dataset in the space of inverse Gaussian Deep FHT ($\{x_{0},\mu\}$). Background colors represent interpolated event times $T(\mathbf{p})$ obtained by inverse distance weighting of uncensored training instances. Points correspond to uncensored test patients and individuals surviving beyond the last uncensored time, colored by their observed event times. In both cases, patients with similar event times cluster in contiguous regions, illustrating that the model encodes similarity in terms of process parameters.
In Fig. 3, a colormap based on $T(\mathbf{p})$ is superimposed on the parameter space of both Lévy and inverse Gaussian models trained on the Framingham dataset, with uncensored test instances and individuals surviving beyond the last uncensored time represented as points colored by their observed event times. The plots show that previously unseen patients are typically assigned to regions with comparable event times, indicating that individuals with similar risk cluster together in the space of process parameters. This proves to be true across all datasets and models (see Appendix C).
We also note that the empirical distribution of event times shown in Fig. 3 matches our theoretical knowledge of process dynamics. Patients with longer survival times cluster in regions with favorable initial conditions (i.e., larger distance from the absorbing barrier) and lower diffusion coefficient or drift; processes with these parameter values typically yield longer first-passage times. Conversely, patients experiencing early failure are assigned higher $D$ or $\mu$ and initial conditions closer to the barrier, consistent with shorter first-passage times.
Feature values distribution.  We further illustrate physics-based interpretability by examining the relationship between clinical features and the learned distribution parameters. For instance, in the inverse Gaussian model applied to the GBSG2 dataset, patients from the test set are mapped into the parameter space $\{x_{0},\mu\}$ and tumor grade is visualized by color gradient (Fig. 4, bottom row). A clear pattern emerges: patients with low-grade tumors (grade 1) occupy regions with favorable initial conditions and weaker drift (low-risk), while those with high-grade tumors (grade 3) cluster near the absorbing barrier with stronger drift (high-risk). This distribution is consistent with established clinical understanding of tumor aggressiveness and prognosis.
A similar analysis for the Lévy model on the Framingham dataset reveals that individuals with higher systolic blood pressure (sbp) and diastolic blood pressure (dbp) are more frequently located in the high-risk region of the $\{x_{0},D\}$ space (Fig. 4, top row). For a given diffusion coefficient $D$, these patients tend to have initial conditions $x_{0}$ closer to the absorbing barrier, linking blood pressure to the only deterministic component of the underlying Brownian motion. This alignment between clinical risk factors and model-derived parameters further supports the physics-based interpretability of DeepFHT.
Final remarks.  We conclude this section by noting a phenomenon visible in Figs. 3,  4: the apparent collinearity between process parameters (or their logarithms). This effect, observed in both Lévy and inverse Gaussian models, leads to data points concentrating along one-dimensional manifolds in parameter space. As reported in the FHT literature [12], this behavior can be attributed to the limited complexity of publicly available datasets rather than to intrinsic shortcomings of the models themselves. We reinforce this notion by showing that the parameter space of the synthetic NonPH dataset does not exhibit this effect (see Appendix C).
Figure 4: Feature–parameter relationships in the Lévy and inverse Gaussian DeepFHT models. Top: Framingham dataset with Lévy model, showing the parameter space $\{x_{0},D\}$ colored by systolic (left) and diastolic (right) blood pressure. Bottom: GBSG2 dataset with inverse Gaussian model, showing the parameter space $\{x_{0},\mu\}$ for patients with grade 1 (left) and grade 3 (right) tumors. In both cases, clinical risk factors align with model-derived high-risk regions (small $x_{0}$, large $D$ or $\mu$), supporting the physical interpretability of the parameterization.
VI Discussion
We introduced DeepFHT, a neural survival model that combines the flexibility of deep learning with the physics-based interpretability of first-passage-time distributions. In this work, we focused on diffusion models, for which closed-form survival functions can be derived and whose parameters admit a clear physical interpretation. Importantly, the mapping from input variables to process parameters can be implemented with arbitrarily complex neural networks, conferring high modeling flexibility and strong fitting power while maintaining physics-based interpretability through the underlying diffusion parameters.
A central advantage of DeepFHT lies in its physics-based interpretability. By mapping individuals to points in parameter space, the model allows us to understand survival in terms of well-characterized process parameters such as initial position, drift, and diffusion. This representation provides intuitive visualizations of survival dynamics and encodes similarity between patient outcomes through distance in parameter space. In this space, DeepFHT reveals clustering of patients by event times and recovers known associations between clinical features and disease outcomes. Such interpretability sidesteps the classical issue of explainability in deep survival models and can be especially valuable in clinical settings where transparency is essential.
Finally, across multiple clinical datasets, DeepFHT achieved predictive performance comparable to or exceeding Cox regression, with clear gains in proof-of-concept scenarios characterized by nonlinear effects and violations of the proportional hazards assumption. Overall, the similar performance of the two approaches suggests limited expressiveness intrinsic to the available datasets. We expect that testing on richer datasets will highlight the value of incorporating deep learning techniques and stochastic models into modern survival analysis.
Appendix A Model architecture and hyperparameters
Hyperparameter configurations  Table 1 presents the best hyperparameter configurations obtained via cross-validation, for each dataset and distribution. For all configurations, batch normalization is applied after each hidden layer for improved training stability. Dropout is subsequently applied, if selected during cross-validation, before the activation function.
The tuned hyperparameters are:
  • Hidden sizes: Model architecture, as a tuple indicating for each hidden layer.
  • Epochs: Number of training epochs.
  • BS: Batch size, computed as $2^{n}$, where $n$ is the original tuned exponent.
  • LR: Learning rate.
  • DO: Dropout probability.
  • Activation: Activation function used in hidden layers.
Table 1: Best hyperparameters per dataset and distribution. Batch normalization is always applied after each hidden layer.
Dataset Model Hidden Sizes Activation DO Epochs BS LR
GBSG2 Levy 16 tanh 0.2 150 64 0.0012
InverseGaussian 16,16 elu 0.5 350 32 0.0005
Framingham Levy 16,16 tanh 0.5 150 64 0.0011
InverseGaussian 16,16 elu 0.2 300 32 0.0003
SUPPORT2 Levy 32,16 elu 0.3 300 32 0.0003
InverseGaussian 32,16 relu 0.4 200 32 0.0001
NonPH Levy 16,16 relu 0.0 450 256 0.0039
InverseGaussian 16,16 relu 0.0 200 64 0.0028
Cross validation scores  In Table 2 we present validation C-index scores of our model across datasets and distributions. For each pair, we report the mean score across cross-validation folds and the corresponding standard deviation.
Model GBSG2 Framingham SUPPORT2 NonPH
Lévy $0.694\pm 0.017$ $0.705\pm 0.012$ $0.854\pm 0.004$ $0.840\pm 0.004$
Inverse Gaussian $0.692\pm 0.024$ $0.700\pm 0.013$ $0.854\pm 0.005$ $0.837\pm 0.010$
Table 2: Cross-validation C-indices for Lévy and inverse Gaussian models across datasets.
Appendix B Synthetic dataset generation
The synthetic dataset NonPH was constructed to strongly violate Cox Regression assumptions, namely proportional hazards and linearity. Here we describe in detail the procedure used to generate it.
First, we built a matrix of 10,000 observations, each with 20 features $x_{1},\ldots,x_{20}$ randomly sampled from a normal distribution. Event times were randomly sampled within the interval $[0,10]$, which was discretized into 1,000 equally sized sub-intervals, plus an additional unbounded $10,+\infty[$ interval. For each observation, we defined a discrete probability distribution over these intervals as the probability that an event occurs in them according to the a specific survival function parametrized by the features. A single interval was then sampled, with its lower bound taken as the event time for bounded intervals or as the censoring time for the unbounded interval. This ensured that all events occurred before time 10 and all censoring at time 10. To enforce consistency, additional censoring was applied so that 25% of observations were censored. Finally, a random subset of 2,400 observations was retained, preserving the censoring ratio, both to reduce computational cost and to reflect a more realistic dataset size.
The survival distribution used for the event sampling was defined as follows. The time interval $[0,10]$ was partitioned into 16 intervals of equal length where the distribution density is a constant on each interval with value $p_{i}$ for $i=1,\ldots,16$. The $p_{i}$ were computed using the soft-max function on the first 16 features (one for each time interval) multiplied by 16: $p_{i}=\frac{\exp(16x_{i})}{\sum^{16}_{j=1}\exp(16x_{j})}$.
The softmax function can be interpreted as a smooth approximation of the argmax operator. When all inputs are scaled by a coefficient $\beta$, the approximation becomes improves: as $\beta$ increases, the probability $p_{i}$ associated with the largest input $x_{i}$ approaches 1, while the others tend toward 0. We set $\beta=16$ (coincidentally matching the number of intervals) so that, for most individuals, the event probability is concentrated in a single time interval. With a smaller value such as $\beta=1$, the resulting distributions would be flatter and therefore easier to approximate for PH assuming methods.
Appendix C Physics-based explainability: time interpolation plots
Figure 5: Time interpolation in parameter space across models for GBSG2, SUPPORT2 and NonPH datasets. Notice the absence of collinearity between parameters for the synthetic NonPH datasets, as discussed in Sec. V.2